OpenStudy (anonymous):

Solve the following separable ODE xy^1 + y = 0, y(4) = 6

OpenStudy (anonymous):

\[xy'+y=0~~\iff~~y'=-\frac{y}{x}~~\implies~~\frac{dy}{y}=-\frac{dx}{x}\]